Fill in the blank. Speed Car Rental company found that the tire lifetime for their vehicles has a mean of 45,964.91 miles and standard deviation of 5,844.84 miles. In a sample of 70 tires tested, there is a 56% chance that the average tire lifetime is less than ________ miles.

Question 3 options:

1)

45,859.44

2)

46,070.38

3)

There is not enough information to determine this.

4)

45,082.52

5)

46,847.3
Answer #1

Given that,

mean = \mu = 45964.91

standard deviation = \sigma = 5844.84

n = 70

\mu\bar x = 45964.91

\sigma\bar x = \sigma / \sqrt n = 5844.84/\sqrt70=698.5920

Using standard normal table,

P(Z < z) =56 %

= P(Z < z) = 0.56

= P(Z <0.151 ) = 0.56

z = 0.151

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

\bar x =0.151 *698.5920 +45964.91

\bar x = 46070.38